Deliora69:

Stacey and Michele both babysit in the evenings. Stacey’s hourly rate is $10 for one child plus $2 per additional child. Michele’s hourly rate is $8 for one child plus $3 per additional child.

supie:

choices?

generation2006:

M=s M=8+3x S=10+2x 10+2x=8+3x x=?

jhonyy9:

ok but what is the question here ?